A company has $1,315 in inventory, $4,746 in net fixed assets, $610 in accounts receivable, $262 in cash, $554 in accounts payable, and $5,341 in equity. What is the company's long-term debt?

  • $1,075

  • $1,217

  • $1,536

  • $1,592

  • $1,038
Answer #1

Total assets=inventory+net fixed assets+AR+Cash

=1,315+4,746+610+262=6933

Total assets=Total liabilities+Total equity

6933=AP+Long term debt+5341

Long term debt=6933-554-5341

=$1038